OpenStudy (anonymous):

find dy/dx by implicit differentiation: x^2[x-y]^2=x^2-y^2

OpenStudy (callisto):

Do you know how to differentiate xy with respect to x?

OpenStudy (anonymous):

yes I just confused after doing the chain rule

OpenStudy (callisto):

Would you mind showing what you have got so far?

OpenStudy (anonymous):

x^2[x-y]^2=x^2-y^2 2x[x-y]^2+x^2(2)(x-y)(1-dy/dx)=2x-2ydy/dx 2x[x-y]^2+x^2(2)(x-y)(1-dy/dx)+2ydy/dx=2x 1-dy/dx+2ydy/dx=2x-2x[x-y]^2-x^2(2)(x-y)

OpenStudy (callisto):

\[x^2(x-y)^2=x^2-y^2\] ^This is the question, right?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

First, let's simplify it a bit. Consider the right side, can you factorize x^2 - y^2?

OpenStudy (anonymous):

the derivative of the right side is 2x-2ydy/dx

OpenStudy (callisto):

Denote dy/dx as y' \[x^2(x-y)^2=x^2-y^2\]Differentiate both sides w.r.t. x \[2x[x-y]^2+x^2(2)(x-y)(1-y')=2x-2yy'\] Expand the term \(2x^2(x-y)(1-y')\) \((2x^3 - 2x^2y)(1-y') = 2x^3-2x^2y - y'(2x^3 - 2x^2y)\) So, we get \[2x[x-y]^2+ 2x^3-2x^2y - y'(2x^3 - 2x^2y)=2x-2yy'\] Put the terms with y' on the same side: \[2x[x-y]^2+ 2x^3-2x^2y - 2x= y'(2x^3 - 2x^2y)-2yy'\] Take out y' as the common factor and isolate it.

OpenStudy (callisto):

Another way I am thinking is: \[x^2[x-y]^2=x^2-y^2\]Factorize the right side, we get \[x^2[x-y]^2=(x-y)(x+y)\] Divide both sides by x-y, for x≠y. Then, we get \[x^2(x-y)=x+y\]Expand the left side \[x^3-x^2y=x+y\]Now we can differentiate the equation.

OpenStudy (anonymous):

I think the your second method is great idea

OpenStudy (anonymous):

a lot faster and for someone who doesn't understand implicit differentiation very well this method allows you to see where it comes into play

OpenStudy (anonymous):

Thank you so much

OpenStudy (callisto):

Usually, it is better if you can simplify the expression a bit :) You're welcome~

OpenStudy (anonymous):

can i ask you another math question

OpenStudy (callisto):

Yes :) But better ask a new question in a new post

OpenStudy (anonymous):

ok thanks

OpenStudy (callisto):

Welcome again